# Real-time Scattering in $\mathbf{\phi^{4}}$ Theory using Matrix Product States

### v1

@@ Abstract
We investigate the critical behavior and real-time scattering dynamics of the interacting $\phi^{4}$ quantum field theory in $(1+1)$ dimensions using uniform matrix product states and the time-dependent variational principle. A finite-entanglement scaling analysis at $\lambda=0.8$ bounds the critical mass-squared to $\mu_{c}^{2}\in[-0.3190,-0.3185]$ and provides a quantitative map of the symmetric, near-critical, weakly broken, and deeply broken regimes. Using these ground states as asymptotic vacua, we simulate two-particle collisions in a sandwich geometry and extract the elastic scattering probability $P_{11\to 11}(E)$ and Wigner time delay $\Delta t(E)$ following the prescription of Jha et al. [1]. We find strongly inelastic scattering in the symmetric phase ($P_{11\to 11}\simeq 0.63$, $\Delta t\simeq-180$ for $\mu^{2}=0.2$), almost perfectly elastic collisions in the spontaneously broken phase ($P_{11\to 11}\simeq 0.998$, $\Delta t\simeq-270$ for $\mu^{2}=-0.2$ and $P_{11\to 11}\simeq 1$, $\Delta t\simeq-177.781$ for $\mu^{2}=-0.5$), and a breakdown of the sandwich evolution precisely at the critical coupling, which provides a dynamical signature of the quantum critical point. These results demonstrate that TDVP-based uniform matrix product states can probe nonperturbative scattering and critical dynamics in lattice $\phi^{4}$ theory with controlled entanglement truncation.
@@ I Introduction
$S_{\phi^{4}}=\int d^{2}x\,\bigg[\frac{1}{2}(\partial_{\nu}\phi)(\partial^{\nu}\phi)-\bigg(\frac{1}{2}\mu^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}\bigg)\bigg],$ (1)
where $\phi(x)$ is a real scalar field, $\mu^{2}$ is the bare mass-squared parameter, and $\lambda>0$ controls the strength of the quartic self-interaction. The $\phi^{4}$ model underlies the field-theoretic description of classical and quantum critical phenomena [7, 8, 9, 10] and provides a canonical realization of Ising universality in $(1+1)$ dimensions [11, 6].
@@
In parallel, there is a growing effort to compute real-time scattering amplitudes directly on quantum computers. Algorithmic proposals for simulating scattering in scalar quantum field theories on fault-tolerant devices have been put forward in Ref. [22], and recent proof-of-principle digital quantum simulations have begun to implement scattering dynamics for scalar and Ising-like field theories using up to $\mathcal{O}(10^{2})$ qubits, W-state encodings, and Hamiltonian truncation on noisy intermediate-scale hardware [23, 24, 25].
@@
In $(1+1)$ dimensions, tensor-network methods — in particular matrix product states (MPS) [26, 27, 28] combined with the time-dependent variational principle (TDVP) and tangent-space methods [29, 30, 31, 32] — provide a complementary route to real-time evolution in the thermodynamic limit. This approach has enabled simulations of false-vacuum collisions, Schwinger pair production, and gauge dynamics [33, 34, 35], and was recently used by Jha et al. to study real-time scattering in Ising field theory [1].
@@
Here we extend TDVP-based uniform MPS (uMPS) methods to the interacting $\phi^{4}$ field theory with the goal of probing both its critical behavior and real-time two-particle scattering across the phase diagram. We focus on a fixed quartic coupling $\lambda=0.8$ and use finite-entanglement scaling (FES) [26] to locate the quantum critical point and characterize the adjacent phases. The FES analysis yields an estimate for the critical mass-squared $\mu_{c}^{2}$ and provides a quantitative map of the symmetric, near-critical, and spontaneously broken regimes that we subsequently explore via scattering simulations.
We discretize the continuum theory with lattice spacing $a$ and work in units $a=\hbar=c=1$. The Hamiltonian is written as $H=\sum_{i}h_{i,i+1}$ with
$h_{\mathrm{i,i+1}}=\frac{1}{2}\left(\pi_{i}^{2}+\pi_{i+1}^{2}\right)+\frac{1}{2}\left(\phi_{i}-\phi_{i+1}\right)^{2}+\frac{1}{2}\mu^{2}\left(\phi_{i}^{2}+\phi_{i+1}^{2}\right)+\frac{\lambda}{4!}\left(\phi_{i}^{4}+\phi_{i+1}^{4}\right),$ (2)
where on-site terms are split evenly across bonds so that $\sum_{i}h_{i,i+1}$ reproduces the standard discretization of $\frac{1}{2}\pi^{2}+\frac{1}{2}(\nabla\phi)^{2}+\frac{1}{2}\mu^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}$. We take $\phi_{i},\pi_{j}]=i\delta_{ij}$, and treat $(\mu,\lambda)$ as bare lattice couplings; when needed we normal-order with respect to a reference mass to control the additive renormalization in (1+1)D.
The many-body state $|\psi\rangle$ is written as a translation-invariant uMPS
$|\psi(A)\rangle=\sum_{\{s_{i}\}}v_{L}^{\dagger}\Bigg(\prod_{n\in\mathbb{Z}}A^{s_{n}}\Bigg)v_{R}\,|\{s_{i}\}\rangle,$ (3)
where $A^{s}$ is a $D\times D$ matrix, identical on every site, $D$ is the bond dimension, and $s$ labels a $d$-dimensional truncated local Hilbert space.
@@
Working directly in the thermodynamic limit, a finite bond dimension $D$ imposes an effective correlation length $\xi_{D}$ even at criticality. Near a conformal fixed point, the half-chain von Neumann entropy $S$ obeys the finite-entanglement scaling relation
@@
with $c$ the central charge and $\xi_{D}\propto D^{\kappa}$. This relation enables us to estimate the critical coupling $\mu_{c}^{2}$ and verify the Ising universality class directly from uMPS ground states, without changing system size. The resulting bracket for $\mu_{c}^{2}$ then serves as input for our real-time scattering simulations in the different phases.
@@ II Methods
Finite-entanglement scaling.— As discussed in the Introduction, a uniform MPS (uMPS) with finite bond dimension $D$ induces an effective correlation length $\xi_{D}$ and the entropy–length relation Eq.(4) near a conformal fixed point. Here we describe how we use this relation in practice to locate the critical coupling and test universality.
@@
For a fixed quartic coupling $\lambda=0.8$ we compute translationally invariant ground states $|\psi_{D}(r)\rangle$ for a set of bond dimensions $\{D\}$ and a sweep of mass parameters $r$ across the putative critical region. For each pair $(r,D)$ we extract the correlation length $\xi_{D}(r)$ from the subleading eigenvalue of the uMPS transfer matrix and obtain the bipartite entropy $S_{D}(r)$ from the Schmidt spectrum. We then fit, at fixed $r$, $S_{D}(r)$ as a function of $\log\xi_{D}(r)$ to a linear form $S_{D}(r)\simeq a(r)\,\log\xi_{D}(r)+b(r),$ and interpret the slope as an effective central charge, $c_{\mathrm{eff}}(r)=6\,a(r)$.
The resulting $c_{\mathrm{eff}}(r)$ profile exhibits a pronounced peak as $r$ approaches criticality. We identify a narrow interval in $r$ that contains this peak and take it as a bracket for the critical coupling,
@@
with the peak position furnishing a central estimate and variations of the fit window and $D$-range providing an uncertainty. As a minimal consistency check we also test data collapse of a local observable (e.g. the order parameter) when plotted as a function of $(r-r_{c})\,\xi_{D}^{1/\nu}$, with $\nu=1$ (Ising). This confirms that the extracted $r_{c}$ is compatible with Ising universality within our accessible bond dimensions. The FES analysis thus yields a quantitatively controlled estimate of the critical mass-squared and a map of the surrounding parameter space that we use to select representative points in the symmetric, near-critical, weakly broken, and deeply broken regimes.
Real-time scattering protocol.— To probe real-time scattering we follow the sandwich geometry introduced in Ref. [1]. For each choice of mass parameter $r$ at fixed quartic coupling $\lambda=0.8$ we take the corresponding uMPS ground state $|\Omega(r)\rangle$ as an approximation to the asymptotic vacuum. We then imprint two opposite-momentum wave packets on top of $|\Omega(r)\rangle$ at well-separated positions,
$|\Psi(0)\rangle=\hat{B}^{\dagger}_{-k}(x_{L})\,\hat{B}^{\dagger}_{+k}(x_{R})\,|\Omega(r)\rangle,$ (6)
where $\hat{B}^{\dagger}_{\pm k}(x)$ creates a localized quasiparticle with central momentum $\pm k$ and a smooth spatial profile such that the initial packets are approximately non-overlapping. The subsequent evolution $|\Psi(t)\rangle$ is obtained by integrating the TDVP equations for a non-uniform MPS defined on a finite window that contains both packets and is embedded in the uniform background $|\Omega(r)\rangle$.
We monitor the local field expectation value $\langle\phi_{n}(t)\rangle$ and its connected correlators throughout the collision. Once the outgoing packets are again well separated, we analyze the asymptotic state to extract the elastic scattering probability $P_{11\to 11}(E)$ at the corresponding center-of-mass energy $E$. In particular, we employ the diagonal projection scheme of Ref. [1], adapted to the $\phi^{4}$ Hamiltonian, which yields a numerically stable estimate of the probability that both outgoing packets contain exactly one quasiparticle. The Wigner time delay $\Delta t(E)$ is obtained from the shift in the peak position of the outgoing wave packets relative to a non-interacting reference evolution at the same energy. This protocol is applied at representative points in the symmetric, near-critical, weakly broken, and deeply broken regimes selected by the FES analysis above.

### v2

@@ Abstract
We investigate the critical behavior and real-time scattering dynamics of the interacting $\phi^{4}$ quantum field theory in (1+1)-dimensions using uniform matrix product states (uMPS) and the time-dependent variational principle (TDVP). A finite-entanglement scaling analysis at $\lambda=0.8$ bounds the critical mass-squared to $\mu_{c}^{2}\in]-0.2595,-0.2594[$ and provides a quantitative map of the symmetric, near-critical, and spontaneously broken regimes. Using these ground states as asymptotic vacua, we simulate two-particle collisions in a sandwich geometry and extract the elastic scattering probability $P_{11\to 11}(E)$ and Wigner time delay $\Delta t(E)$ using a sandwich geometry protocol. We find strongly inelastic scattering in the symmetric phase ($P_{11\to 11}\simeq 0.712$, $\Delta t\simeq-158$ for $\mu^{2}=+0.2$) and almost perfectly elastic collisions in the spontaneously broken phase ($P_{11\to 11}\simeq 1$, $\Delta t\simeq-108$ for $\mu^{2}=-0.1$ and $P_{11\to 11}\simeq 1$, $\Delta t\simeq-177.781$ for $\mu^{2}=-0.5$). Crucially, the scattering protocol exhibits a distinctive divergence near the critical coupling; we show that this behavior serves as a dynamical signature of the quantum critical point, arising directly from the closing of the mass gap. These results demonstrate that TDVP-based uMPS can effectively probe nonperturbative scattering and critical dynamics in lattice field theories with controlled entanglement truncation.
@@ I Introduction
$\mathcal{S}_{\phi^{4}}$ $=\int d^{2}x\,\mathcal{L}_{\phi^{4}}$
$=\int d^{2}x\,\bigg[\frac{1}{2}(\partial_{0}\phi)(\partial^{0}\phi)-\frac{1}{2}(\partial_{x}\phi)(\partial^{x}\phi)$
$~~~~~~~~~~~~~~~~~~~~~~~-\bigg(\frac{1}{2}m^{2}\phi^{2}+\frac{g}{4!}\phi^{4}\bigg)\bigg],$ (1)
where $\mathcal{L}_{\phi^{4}}$ is the Lagrangian density, $\phi\equiv\phi(x)$ is a real scalar field, $m$ is the mass parameter, and $g>0$ controls the strength of the quartic self-interaction. The $\phi^{4}$ model underlies the field-theoretic description of classical and quantum critical phenomena [6, 7, 8, 9] and provides a canonical realization of Ising universality in $(1+1)$ dimensions [10, 5].
@@
In parallel, there is a growing effort to compute real-time scattering amplitudes directly on quantum computers. Algorithmic proposals for simulating scattering in scalar quantum field theories on fault-tolerant devices have been put forward in Ref. [21], and recent proof-of-principle digital quantum simulations have begun to implement scattering dynamics for scalar and Ising-like field theories using up to $\mathcal{O}(10^{2})$ qubits, W-state encodings, and Hamiltonian truncation on noisy intermediate-scale hardware [22, 23, 24]. While Hamiltonian truncation has successfully established high-precision benchmarks for the static phase diagram [25], tensor network methods provide a more direct route to simulating real-time dynamics in the thermodynamic limit.
@@
In $(1+1)$ dimensions, matrix product states (MPS) [26, 27, 28], combined with the time-dependent variational principle (TDVP) and tangent-space methods [29, 30, 31, 32], provide a complementary route to real-time evolution in the thermodynamic limit. This approach has enabled simulations of false-vacuum collisions, Schwinger pair production, and gauge dynamics [33, 34, 35], and was recently used by Jha et al. to study real-time scattering in Ising field theory [36].
@@
Here we extend TDVP-based uniform MPS (uMPS) methods to the interacting $\phi^{4}$ field theory with the goal of probing both its critical behavior and real-time two-particle scattering across the phase diagram. We focus on a fixed self-interaction parameter and use finite-entanglement scaling (FES) [26] to locate the quantum critical point and characterize the adjacent phases. The FES analysis yields an estimate for the critical mass and provides a quantitative map of the symmetric, near-critical, and spontaneously broken regimes that we subsequently explore via scattering simulations.
For our computations, we require a discretized Hamiltonian. Therefore, starting with the Legendre transformation of the Lagrangian density $\mathcal{L}_{\phi^{4}}$ introduced in Eq.(1), we discretize the spatial dimension onto a one-dimensional lattice with spacing $a$. By approximating the spatial derivative using a finite difference, we obtain the following Hamiltonian:
$H_{\phi^{4}}=\frac{1}{a}\sum_{j}\left(\frac{1}{2}\pi_{j}^{2}+\frac{1}{2}\left(\phi_{j+1}-\phi_{j}\right)^{2}+\frac{1}{2}\mu^{2}\phi_{j}^{2}+\frac{\lambda}{4!}\phi_{j}^{4}\right),$ (2)
with $\pi_{j}$ representing the conjugate momenta, $\mu^{2}\equiv a^{2}m^{2}$ the bare mass-squared parameter, and $\lambda\equiv a^{2}g$ the quartic coupling. We can neglect the overall factor of $\frac{1}{a}$ and in our work we have set $a=c=\hbar=1$. The system respects the canonical commutation relation $\phi_{j},\pi_{k}]=i\delta_{jk}$ and treats ($\mu,\lambda$) as bare lattice couplings. Local Hilbert space dimensions are truncated to a local value $d$.
We renormalize the Hamiltonian by subtracting the local ground-state energy density to ensure finite total energy. As noted by Sugihara [37], standard normal ordering is defined in the Fock basis and is non-local in a real-space lattice representation. We therefore employ a local subtraction scheme consistent with the vacuum energy removal used in Hamiltonian truncation estimates.
To describe the state of the system we choose the MPS representation. In this framework, the lattice chain is represented by sites that carry information about the system via square $D\times D$ matrices, $A^{s_{n}}$, where $D$ is the maximum bond dimension of the MPS. This parameter controls the accuracy of the ansatz and sets an upper bound on the entanglement entropy, $S\leq\log(D)$.
MPS can generically take the form
$|\Psi\rangle=\sum_{\{s\}}\left(\vec{v}_{L}^{\dagger}\dots A_{-1}^{(s-1)}A_{0}^{(s_{0})}A_{1}^{(s_{1})}\dots\vec{v}_{R}\right)|s\rangle,$ (3)
where $|s\rangle=|\dots s_{-1}s_{0}s_{1}\dots\rangle$, and $\vec{v}_{L}$ and $\vec{v}_{R}$ are vectors that live at spatial infinity and are irrelevant in the thermodynamic limit.
One can see that Eq.(3) is invariant under gauge transformations of the form $A_{n}^{(s_{n})}\rightarrow gA_{n}^{(s_{n})}g^{-1}$; in particular, there exists a gauge transformation that allows all matrices to become site-independent, i.e. $A^{s_{n}}_{n}=A^{s_{n}}$ for all sites $s_{n}$. In this case, the representation is known as a uniform matrix product state (uMPS) and takes the form
$|\Psi\rangle=\sum_{\{s_{i}\}}\vec{v}_{L}^{\dagger}\left(\prod_{n\in\mathbb{Z}}A^{(s_{n})}\right)\vec{v}_{R}\;|s\rangle$ (4)
The central object in MPS is the transfer matrix,
$E=\sum_{s=1}^{d}A^{s}\otimes\bar{A}^{s},$ (5)
a completely positive map whose determinant eigenvalue can be scaled to 1. In the generic case the leading eigenvalue is non-degenerate with positive left and right fixed points $l$ and $r$, normalized by $\mathrm{Tr}(lr)=1$. These properties ensure well-defined thermodynamic quantities and independence from boundary vectors.
Local observables in this representation are evaluated by contracting transfer matrix fixed points. For some operator $O$, with $\langle O\rangle=\langle l|E_{O}|r\rangle$,
$E_{O}=\sum_{s,t}\langle s|O|t\rangle A^{s}\otimes\bar{A}^{t}.$ (6)
For a two-site operator, say the nearest-neighbor Hamiltonian $\langle h\rangle=\langle l|E_{h}|r\rangle$,
$E_{h}=\sum_{s_{1},s_{2},t_{1},t_{2}}\langle s_{1}s_{2}|h|t_{1}t_{2}\rangle\left(A^{s_{1}}A^{s_{2}}\right)\otimes\left(\bar{A}^{t_{1}}\bar{A}^{t_{2}}\right).$ (7)
For one-dimensional (1D) systems, the matrix product state (MPS) ansatz serves as a high-fidelity variational manifold $\mathcal{M}_{\text{MPS}}$ that efficiently captures the required area law entanglement structure of ground and low-energy excited states. The time-dependent variational principle (TDVP) [31, 32, 29] provides a rigorously grounded framework for performing both real- and imaginary-time evolution while strictly confining the state’s trajectory to this low-dimensional manifold, $|{\Psi(A(t))\rangle}\in\mathcal{M}_{\text{MPS}}$.
Figure 1: Time-Dependent Variational Principle as depicted in Haegeman et al. [29]
The TDVP procedure commences with the time-dependent Schrödinger equation,
$i\frac{\partial}{\partial t}|{\Psi(t)}\rangle=\hat{H}|{\Psi(t)}\rangle.$ (8)
While the exact evolution vector $\hat{H}|{\Psi(A)}\rangle$ generally points out of the manifold $\mathcal{M}_{\text{MPS}}$, the TDVP seeks an optimal path $|{\Psi(A(t))}\rangle$ by projecting the instantaneous evolution onto the tangent space $T_{A}\mathcal{M}_{\text{MPS}}$ at the current state, $|{\Psi(A)}\rangle$. This optimization is formulated as a minimization problem for the residual error vector:
[…]
where $|{\Phi(B;A)}\rangle=B^{i}\frac{\partial}{\partial A^{i}}|{\Psi(A)}\rangle=B^{i}|{\partial_{i}\Psi(A)}\rangle$ is a generic tangent vector parametrized by the tensor $B$, and $A$ denotes the collective MPS variational parameters.
The minimization condition is mathematically equivalent to demanding that the residual error vector be orthogonal to the tangent space, which yields the TDVP equations of motion:
$\langle{\partial_{\bar{j}}\Psi|\partial_{i}\Psi}\dot{A}^{i}\rangle=-i\langle{\partial_{\bar{j}}\Psi|\hat{H}|\Psi}\rangle.$ (10)
The TDVP evolution is thus an orthogonal projection of the exact dynamics onto the MPS manifold (see Fig. 1), represented by the flow equation:
$i\frac{\partial}{\partial t}|{\Psi(A(t))}\rangle=P_{A(t)}\hat{H}|{\Psi(A(t))}\rangle,$ (11)
where $P_{A}$ is the orthogonal projector onto $T_{A}\mathcal{M}_{\text{MPS}}$.
@@
In this work we first leverage the scaling of entropy with bond dimension, noting that a finite $D$ imposes an effective correlation length $\xi_{D}$ even at criticality. Near a conformal fixed point, the half-chain von Neumann entropy $S$ obeys the finite-entanglement scaling relation
@@
with $c$ the central charge and $\xi_{D}\propto D^{\kappa}$. This relation enables us to estimate the critical coupling $\mu^{2}_{c}$ and verify the Ising universality class directly from uMPS ground states, without changing system size. The resulting bracket for $\mu_{c}^{2}$ then serves as input for our real-time scattering simulations in the different phases.
@@ II Methods
II.1 Finite-entanglement scaling
As discussed in the Introduction, a uniform MPS (uMPS) with finite bond dimension $D$ induces an effective correlation length $\xi_{D}$ and the entropy–length relation Eq.(12) near a conformal fixed point. Here we describe how we use this relation in practice to locate the critical coupling and test universality.
@@ II.1 Finite-entanglement scaling
For a fixed quartic coupling $\lambda$ we compute translationally invariant ground states $|\psi_{D}(r)\rangle$ for a set of bond dimensions $\{D\}$ and a sweep of mass parameters $r$ across the critical region. For each pair $(r,D)$ we extract the correlation length $\xi_{D}(r)$ from the subleading eigenvalue of the uMPS transfer matrix and obtain the bipartite entropy $S_{D}(r)$ from the Schmidt spectrum. We then fit, at fixed $r$, $S_{D}(r)$ as a function of $\log\xi_{D}(r)$ to a linear form $S_{D}(r)\simeq a(r)\,\log\xi_{D}(r)+b(r),$ and interpret the slope as an effective central charge, $c_{\mathrm{eff}}(r)=6\,a(r)$.
The resulting $c_{\mathrm{eff}}(r)$ profile exhibits a convergence to $c_{\mathrm{eff}}=0.5$ as $r$ approaches criticality. We identify a narrow interval in $r$ that contains this value and take it as a bracket for the critical coupling,
@@
with the peak position yielding a central estimate and variations of the fit window and $D$-range providing an uncertainty. As a complementary consistency check we also test data collapse of a local observable (e.g. the order parameter) when plotted as a function of $(r-r_{c})\,\xi_{D}^{1/\nu}$, with $\nu=1$ (Ising). This confirms that the extracted $r_{c}$ is compatible with Ising universality within our accessible bond dimensions. The FES analysis thus yields a quantitatively controlled estimate of the critical mass-squared and a map of the surrounding parameter space that we use to select representative points in the symmetric, near-critical, weakly broken, and deeply broken regimes.
II.2 Real-time scattering protocol
To probe real-time scattering we follow the sandwich geometry introduced in Ref. [36]. For each choice of mass parameter $r$ at a fixed quartic coupling $\lambda$, a fixed maximal bond dimension $D$, and a fixed Hilbert space truncation $d$ we take the corresponding uMPS ground state as an approximation to the asymptotic vacuum. Starting from said uMPS ground state, we create a non-uniform ‘sandwich’ MPS of length $N$ embedded in that uniform background, see Fig. 2. This gives a window where we can imprint and evolve localized excitations while the far left/right halves remain vacuum-like.
To build these clean quasiparticle wavepackets, we construct a tangent-space vector $B$ from the local physical operator $\phi$ by projecting $O\cdot A$ onto the left-gauge orthogonal complement of the uMPS manifold:
$V^{s}=\sum_{s^{\prime}}O_{ss^{\prime}}A^{s^{\prime}}\;,~~~B=V-A\cdot\left(\sum_{s}A^{s\dagger}V^{s}\right),$ (14)
then normalize $B$. This enforces $\sum_{s}A^{s\dagger}B^{s}=0$ so the packet has no vacuum component in the chosen gauge.
Figure 2: Sandwiching B-Tensor Inside a Uniform Ground State
We imprint a packet by adding small copies of $B$ across the window with a Gaussian envelope around a center $n_{0}$ and a plane-wave phase $e^{i\kappa(n-n_{0})}$. This is done gently (small amplitude), in tiles, restoring canonical form and backing off if a local update fails—so the state stays well-conditioned. We do this twice to produce two-counter propagating packets with $\pm\kappa$.
In short, we build $B=P_{\perp}[\phi\cdot A]$ and then imprint Gaussian $\pm\kappa$ packets.
We monitor the local field expectation value $\langle\phi_{n}(t)\rangle$ and its connected correlators throughout the collision.
To extract scattering data, we use a script that projects the evolving state onto a compact “two-packet” subspace and reads off a set of diagonal Fourier amplitudes $A_{k,-k}(t)$ as mentioned in [36]. Practically, it focuses on lattice sites within Gaussian windows around the instantaneous centers of the left- and right-moving packets and enforces a minimum separation $\Delta n$ so that only well-separated pairs are considered. Using a reusable reference sandwich initially in the vacuum-like state, it inserts tiny perturbations $\varepsilon B$ at selected $(i_{L},i_{R})$ site pairs to define basis vectors for this two-$B$ subspace. The algorithm then computes overlaps with the current state, multiplies by the plane-wave phases to target momenta (k,k)(k,-k), accumulates the results across the chosen pairs, and finally normalizes by $\varepsilon^{2}$ to stay in a linear-response regime. To keep the cost and variance under control, pairs with negligible Gaussian weight are dropped and the total pair count is capped. We mainly are concerned with time delay and probability of elastic scattering.
The time delay measures the change in arrival (or departure) time of scattered wave packets compared with non-interacting particles. Physically, it represents how long the interacting particles “linger” in the scattering region because of interaction effects encoded in the phase of the elastic S-matrix element $S_{11\rightarrow 11}(E)$. A positive time delay indicates that the particles spend more time interacting (as if temporarily trapped, e.g., near a resonance), while a negative delay means they pass through faster than in free motion.
We calculate here the Wigner time delay, given by:
$\Delta t=-i\partial_{E}\log{S_{11\rightarrow 11}(E)}.$ (15)
To do so from overlaps at momentum $\kappa_{0}$, we consider nearby momenta $\kappa_{\pm}=\kappa_{0}\pm\delta\kappa$ and compute:
$\phi(t)$ $=\arg\left(\frac{\langle{\kappa_{+},-\kappa_{+}|\psi(t)}\rangle}{\langle{\kappa_{+},-\kappa_{+}|\psi(0)}\rangle}\frac{\langle{\kappa_{-},-\kappa_{-}|\psi(0)}\rangle}{\langle{\kappa_{-},-\kappa_{-}|\psi(t)}\rangle}\right)$
$=2(E_{\kappa_{+}}-E_{\kappa_{-}})t+\Delta(\kappa_{+})+\Delta(\kappa_{-}),$ (16)
this quantity captures the relative phase-shift between components of $|{\psi}\rangle$ with momenta $\kappa_{\pm}$. The first term in $\phi(t)$, due to an expected phase shift from the free theory, is subtracted, and $\Delta(\kappa_{\pm})$ is expanded to first order around $\kappa_{0}$:
$\frac{\Delta(\kappa_{+})-\Delta(\kappa_{-})}{2(E_{\kappa_{+}}-E_{\kappa_{-}})}$ $\approx\partial_{E}\Delta(E)$
$=-i\partial_{E}\log{S_{11\rightarrow 11}(E)}.$ (17)
As a simple estimate of the elastic scattering $1+1\rightarrow 1+1$ probability, the script compares diagonal spectral weight before and after the collision within a narrow band around $\kappa_{0}$. It sums $|A_{\kappa,-\kappa}(t)|^{2}$ over the sampled $\kappa$ values at an early time and at a late time and takes their ratio as $P$
$P_{11\rightarrow 11}=\frac{|A_{\kappa,-\kappa}(t_{f})|^{2}}{|A_{\kappa,-\kappa}(t_{i})|^{2}}.$ (18)
We practically calculate the probability of finding particles of the ‘same kind’ in the $t_{f}$ state as those in the $t_{i}$ state.
This “diagonal-only” approximation is intentionally lightweight: it avoids the expense of reconstructing the full two-momentum distribution $|A_{\kappa,\kappa^{\prime}}|^{2}$ while still providing a stable, interpretable signal for predominantly elastic scattering. The trade-off is that genuinely off-diagonal contributions with $\kappa^{\prime}\neq\kappa$ are ignored, so the estimate is most reliable when the packets are narrow in $\kappa$ and the scattering remains nearly elastic.